OpenStudy (magbak):

What are the foci of the ellipse given by the equation 4x2 + 9y2 – 32x + 18y + 37 = 0? (4, –1 ±square root of 5) (4, –1 ±square root of 13) (4 ±square root of 5 , –1) (4 ±square root of 5 , 1)

OpenStudy (magbak):

@satellite73

OpenStudy (anonymous):

this is a pain in the neck you have to make it look like \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] by completing the square for both the \(x\) and\(y\) terms any idea how to do that?

OpenStudy (magbak):

Yes

OpenStudy (anonymous):

ok let me know what you get when it is in the form \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] and then we can answer the question

OpenStudy (magbak):

Oh no that I do not know.

OpenStudy (anonymous):

well then you can't solve the problem!

OpenStudy (anonymous):

i can walk you through it if you like

OpenStudy (magbak):

yes please

OpenStudy (anonymous):

\[4x^2 + 9y^2 – 32x + 18y + 37 = 0\] group the terms together and get \[4x^2-32x+9y^2+18y=-37\] then factor out the coefficient of the \(x^2\) and \(y^2\) terms go get \[4(x^2-8x)+9(y^2+2y)=-37\] so far so good?

OpenStudy (magbak):

yes

OpenStudy (anonymous):

ok good what is half of 8 ?

OpenStudy (magbak):

4

OpenStudy (magbak):

so it is either a or b

OpenStudy (anonymous):

hold on what is \(4^2\)

OpenStudy (magbak):

16

OpenStudy (anonymous):

ok and finally, what is \(4\times 16\)?

OpenStudy (magbak):

64

OpenStudy (anonymous):

ok so now we can turn \[4(x^2-8x)+9(y^2+2y)=-37\] in to \[4(x-4)^2+9(y^2+2y)=-37+64=27\]

OpenStudy (magbak):

So is it B

OpenStudy (anonymous):

now we have to repeat the process with the \(y\) terms half of 2 is 1 1 squared is 1, 9 times 1 is 9

OpenStudy (magbak):

I think it is B right. /-\

OpenStudy (anonymous):

\[4(x-4)^2+9(y^2+2y)=-37+64=27\] is now \[4(x-4)^2+9(y+1)^2=27+9=36\]

OpenStudy (anonymous):

now divide all by \(36\)

OpenStudy (magbak):

Oh so it is A

OpenStudy (magbak):

Sorry it is because I have been studying since 3PM until now.

OpenStudy (anonymous):

you are guessing, and still not done for sure

OpenStudy (anonymous):

you get \[\frac{(x-4)^2}{9}+\frac{(y+1)^2}{4}=1\]

OpenStudy (magbak):

Oh so it is D now that I understand.

OpenStudy (magbak):

One more question please I am so tiered but I must complete this last question.

OpenStudy (anonymous):

no, it is not D

OpenStudy (magbak):

Oh so what is it?I am so sorry again.

OpenStudy (anonymous):

center is \((4,-1)\) right?

OpenStudy (magbak):

As I said before if it is not written in the question then it is not their.

OpenStudy (anonymous):

and here \(a^2=9\) making \(a=3\), \(b^2=4\) making \(b=2\) and the focus is \(\sqrt{9-4}=\sqrt{5}\) units to the left and right of the center

OpenStudy (magbak):

So it is A as I said previously before.

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

you keep missing C it is always C

OpenStudy (anonymous):

\(\sqrt{5}\) units left of \((4,-1)\) is \((4-\sqrt{5},-1)\)

OpenStudy (magbak):

HA I always guess c but it has been failing me lately.

OpenStudy (magbak):

Thnx JUST ONE MORE AND I WILL STOP DISTRUBING YOU AND I WILL BE OUT OF YOUR LIFE FOR TODAY

OpenStudy (anonymous):

it is a lot of work okay

OpenStudy (magbak):

Also why can I not message you.

OpenStudy (anonymous):

because i have that annoying feature turned off

OpenStudy (magbak):

Oh so can I message you or will I be annoying.

OpenStudy (anonymous):

no one can

OpenStudy (magbak):

I posted the new question. Also ok thanx. Well except for the people that you fanned.